The average daily production of a certain product in a plant is recorded. Registration was carried out for 54 days and an average of 1,350 liters with a standard deviation of 260 liters was obtained. Test the hypothesis at a significance level of 5%, that the average daily production is 1400 liters. What is the p-value of the test?

A) 0.0793

B) 0.0882

C) 0.0623

D) 0.0551
Answer #1

Solution :

This is the left tailed test,  

The null and alternative hypothesis is ,

H0 :  \mu = 1400

Ha : \mu < 1400

Test statistic = z

= (T - \mu ) / \sigma / \sqrt n

= (1350 - 1400) / 260 / \sqrt 54

Test statistic = z = -1.41

P(Z < -1.41)

P-value = 0.0793

correct option is = A